OpenStudy (anonymous):

Find the limit of the function algebraically. -6 -3 Does not exist 0 Im kind of confused on how to do this exactly. Could someone maybe walk me through the steps?

OpenStudy (anonymous):

Oh hold on you cant see the equation

OpenStudy (anonymous):

lim of x approaching -3, (x^2-9)/(x^3+3)

OpenStudy (anonymous):

where you see an x, put a 3 that is all

OpenStudy (anonymous):

negative 3 or just 3

OpenStudy (anonymous):

if you get \(\frac{0}{0}\) then you have more work to do oh right, -3 what you said

OpenStudy (anonymous):

but in this case you get 0 over something that is not zero, answer is just zero

OpenStudy (anonymous):

So i did that and got -18/-24

OpenStudy (anonymous):

you dont get 0 on top because its -3^2-9 which would be -18

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\((-3)^2=9 \) not \(-3^2=-9\)

OpenStudy (anonymous):

OHH my gosh. I keep making such dumb mistakes. duhhh. thank you!!

OpenStudy (anonymous):

lol yw

OpenStudy (anonymous):

Im having trouble with limits. Would you like to help me?

OpenStudy (anonymous):

here are the possible answers The limit does not exist. -1 -4 7

OpenStudy (anonymous):

I graphed it but how do i know when the limit exists.? Do they need to have the same point when they approach -4?

OpenStudy (anonymous):

limit from the left is not equal to the limit from the right, so it does not exist

OpenStudy (anonymous):

yeah right, they need to be teh same

OpenStudy (anonymous):

Okay, thank you tons!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

tons